# Evaluating Dialect Robustness of Language Models via Conversation Understanding

### v2

@@ Abstract
With an evergrowing number of LLMs reporting superlative performance for English, their ability to perform equitably for different dialects of English (i.e., dialect robustness) needs to be ascertained. Specifically, we use English language (US English or Indian English) conversations between humans who play the word-guessing game of ‘taboo‘. We formulate two evaluative tasks: target word prediction (TWP) (i.e., predict the masked target word in a conversation) and target word selection (TWS) (i.e., select the most likely masked target word in a conversation, from among a set of candidate words). Extending MD3, an existing dialectic dataset of taboo-playing conversations, we introduce M-MD3, a target-word-masked version of MD3 with the en-US and en-IN subsets. We create two subsets: en-MV (where en-US is transformed to include dialectal information) and en-TR (where dialectal information is removed from en-IN). We evaluate one open-source (Llama3) and two closed-source (GPT-4/3.5) LLMs. LLMs perform significantly better for US English than Indian English for both TWP and TWS tasks, for all settings, exhibiting marginalisation against the Indian dialect of English. While GPT-based models perform the best, the comparatively smaller models work more equitably after fine-tuning. Our error analysis shows that the LLMs can understand the dialect better after fine-tuning using dialectal data. Our evaluation methodology exhibits a novel way to examine attributes of language models using pre-existing dialogue datasets.
@@ 1 Introduction
refer to Indian English and US English respectively.Large language models (LLMs)111We use ‘language models’ and ‘large language models/LLMs’ interchangeably in this paper. based on Transformers (Vaswani et al. 2017) are the state-of-the-art in natural language processing (NLP), often reporting superlative performance on several NLP tasks (Zhao et al. 2023). These models predominantly use English language data in their pre-training corpus. However, being a widely spoken language, English takes multiple forms in different parts of the world. These forms, called dialects or national varieties of English, collectively constitute the World Englishes (Bolton 2012). While research papers introducing LLMs report performance on English language datasets, recent works highlight the performance gap between US English and other dialects of English for several natural language processing tasks (Joshi et al. 2024).
Our paper follows this line of work of evaluating LLMs for dialects of English via conversation understanding. The choice of conversation understanding as a domain for evaluation emerges from the fact that dialectal features are most visible in free-flowing conversations (Negro and Vietti 2006). Therefore, we investigate the research question:
@@
To address the research question, we use the MD3 (Eisenstein et al. 2023) that consists of manually transcribed dialogues between pairs of human participants where each pair speaks either Indian English or US English. The participants engage in a focused conversation: they play the word-guessing game based on the game of ‘Taboo’ (Wikipedia 2023). In the game, a describer must get a guesser to identify a target word but must not use a set of related words known as restricted words while describing the target word. Using this dataset of dialectal dialogues, we introduce two tasks to evaluate the dialect-robustness of LLMs to understand conversations. They are: (a) Given an input conversation with the target word masked, can the LLM predict the target word? (referred to as target word prediction) (b) Given an input conversation with the target word masked along with a set of candidate target words, can the LLM select the correct target word? (referred to as target word selection). Our approach of masking the target word is similar to Dey and Desarkar (2023), who show that masked word prediction may correlate with automatic dialogue evaluation metrics. Figure 1 shows an example of the two tasks, where the language model predicts ‘Justin Bieber’ for target word prediction, and selects ‘microphone’ among the set of options for target word selection222We run experiments on both the tasks for both US and Indian English conversations. While the examples show expected output, the LLM may or may not produce the same in the case of our experiments. That is the crux of the evaluation.. For the two tasks, we extend MD3 to create a target-word-Masked Multi-Dialect Dataset of Dialogues (M-MD3)333M-MD3 dataset and the related code will be made publicly available at https://github.com/dipankarsrirag/eval-dialect-robust.. M-MD3 consists of (a) conversations between Indian English speakers (en-IN), and conversations between US English speakers (en-US), (b) en-US conversations transformed into en-IN using rule-based perturbations (en-MV), (c) en-IN with dialectal information removed (en-TR). We evaluate the performance of three SOTA large language models (LLMs), one open-source and two closed-source, employing zero-shot prompting on both pre-trained and fine-tuned models (where available).
Our evaluation methodology derives from past work that evaluates LLMs by providing a set of task-specific examples (Wang et al. 2023). Of particular relevance is the work by Chalamalasetti et al. (2023), who generate word game conversations using LLMs and evaluate their ability to predict the target word. The contributions of our work are:
@@
Our evaluation demonstrates a degraded performance in the case of Indian English as compared to US English for all models, supporting existing social disparities between US and Indian culture in the LLM representations (Khandelwal et al. 2024).
@@
Since several LLMs have been deployed as publicly available dialogue agents444ChatGPT https://chat.openai.com/; Accessed on 9th April 2024., it is imperative that they are able to understand conversations for users belonging to diverse English-speaking subgroups. In the case of our paper, this refers to dialectal variations. The rest of the paper is organized as follows. Section 2 introduces our evaluation methodology. The experiment setup and results are in Sections 3 and 4 respectively.
@@ 2 Methodology
We present our method step-by-step, with a detailed overview of our evaluation methodology described in Figure 2. We select two subsets available in MD3: en-IN and en-US, and filter out the conversations where the guesser could not identify the target word. We extend MD3 to include two additional sets of conversations—en-MV and en-TR, and mask the target words in all four subsets to create M-MD3. We ensure that the mask token always appears at the end of the conversation, warranting the use of auto-regressive models. This is done by pruning the conversation to the turn where the guesser utters the target word555Details on the masking method with examples are provided in the Dataset Construction section of the Supplementary material..
@@ 2.1 Extending MD3
Transforming text in en-US to other dialectal English text has been explored for low-resource settings (Held, Ziems, and Yang 2023; Xiao et al. 2023; Liu, Held, and Yang 2023). To evaluate the efficacy of synthetically transformed dialogues, we extend the dataset of dialectal dialogues to include two additional sets of conversations– en-MV and en-TR.
en-MV
We use Multi-VALUE (Ziems et al. 2023) to transform en-US conversations into en-IN conversations. We call this set of conversations created by rule-based transformations en-MV.
en-TR
We prompt GPT-4 Turbo Preview(gpt-4; OpenAI 2024) to remove dialectal information from en-IN. The resultant set of conversations is known as en-TR. The prompt used to generate such conversations is given below:
The use of gpt-4 to transform en-IN conversations sometimes leads to the generation of conversation summaries rather than transformed conversations666More details with examples are discussed in the Transformation Issues section of the Supplementary material.. Due to the varying lengths of speaker turns, transforming en-US conversations using Multi-VALUE occasionally fails to output a result. Such failed transformations are excluded from both the subsets of transformed (en-MV, en-TR) conversations, leading to fewer conversations in en-MV and en-TR as compared to en-US and en-IN, respectively, as shown in Tables 1 and 2.
@@ 2.2 Analysis
We notice a higher number of average turns and words spoken in en-IN conversations compared to en-US conversations. This is due to the en-US speakers being more familiar with the target word compared to en-IN speakers, leading to shorter gameplay time (Eisenstein et al. 2023). The trend is also carried over to the transformed conversations in en-MV (derived from en-US) and en-TR (derived from en-IN).
@@ 2.3 Task Definition
As shown in Figure 3, we mask the target word in the conversations from all four subsets. The target word occurs in the last dialogue turn of the conversation, which is spoken by the guesser777This always holds because of the way we process the conversations.. As a result, we formulate two tasks where the expected output is to fill the correct word at the masked position:
@@
We then use prompting on three LLMs to perform both tasks (TWP and TWS). As LLMs, we choose models that have been optimised to follow natural language instructions. In our case, the instruction is to either predict the masked target word or select a word from candidate words. Specifically, we use one open-source model, namely, Llama 3 70B Chat (llama-3; Llama Team 2024), and two closed-source models, namely, gpt-4 and GPT-3.5 Turbo 0125 (gpt-3.5; Ouyang et al. 2022).
@@ 3 Experiment Setup
We report the performance on pre-trained and fine-tuned versions of LLMs using zero-shot prompting. Fine-tuning is always done ‘in-dialect’ in our case, although there is no reason to believe that cross-dialect fine-tuning is not possible.
@@ 3.1 Model Parameters
Experiments on gpt-4 and gpt-3.5 are conducted using OpenAI’s API888OpenAI API https://platform.openai.com/docs/api-reference; Accessed on 18th April 2024.. gpt-3.5 is fine-tuned for 5 epochs, separately for every subset. We select top_p as 0.2 to restrict variability in output generation.
llama-3 is fine-tuned for 20 epochs, with a batch size of 16, Paged 8-bit AdamW (Dettmers et al. 2022) as the optimiser and a learning rate of 2e-4. We use QLoRA adaptors, targeting all linear layers, as recommended by Dettmers et al. (2023). All experiments on llama-3 were performed using two A100 GPUs.
@@ 3.2 Metrics
We report our results on two metrics: accuracy and similarity. Accuracy is the proportion of conversations where the LLM generated the correct target word. This is a strict metric in that it requires the LLM to generate an exact match to the reference target word. In the case of TWP, the LLM will choose from all the words within its vocabulary, while in the case of TWS, the LLM will choose from the set of candidate target words. Therefore, it is trivial that the accuracy for TWS is expected to be higher than that for TWP. Accuracy metric penalizes models even if the generated target word partially matches with the reference target word in case of multi-word reference target as described in Section 2.2. As similarity, we report the cosine similarity between the Sentence-BERT embeddings (Reimers and Gurevych 2019) of the reference target word and the generated target word. This allows for similar but inexact words generated by the LLM to be acceptable to the similarity score.
@@ 4.1 Quantitative Results
In the case of en-MV, the task performances are consistently lower compared to en-US. For example, in the case of gpt-3.5 and TWS, the similarity scores on the fine-tuned model are 80.8 for en-US and 71.5 for en-MV. This degraded performance shows that the rule-based transformation into en-IN from en-US reduced the understanding capacity of LLMs for the resulting conversations, further strengthening our hypothesis that LLMs perform well for US English dialects compared to any other varieties, similar to findings of Ryan, Held, and Yang (2024).
@@ 4.2 Error Analysis
From Test set of each conversation subset, we randomly select 30 conversations that are mislabeled by gpt-4 and llama-3, and manually analyse errors among all model variants across all subsets of conversations. We summarise the six error categories999Additonal examples for each error category are in the Errors section of the Supplementary material. in Table 4. The error types are:
@@ 5 Related Work
Research in dialect robustness stems from the need for language technologies to be equitable and not reinforce any negative sentiments against a specific linguistic subgroup (Blodgett et al. 2020). LLMs perform poorly on several downstream tasks (such as the tasks in the GLUE benchmark) involving dialects other than mainstream US English (Joshi et al. 2024; Faisal et al. 2024).
Similar to our work, the evaluation of language understanding ability of LLMs has been explored using typical conversation understanding tasks (Chen et al. 2022) like conversation summarisation (Gliwa et al. 2019; Chen et al. 2021), conversation completion (Sai et al. 2020; Ueyama and Kano 2023), or NLU tasks (Faisal et al. 2024). Other approaches involve conversation-based question-answering tasks that also evaluate the reasoning abilities of LLMs (Sun et al. 2019; Qin et al. 2021). Tasks like mask-filling were used to evaluate LLM-generated responses, more specifically Dey and Desarkar (2023) do so by making RoBERTa predict masked keyword utterances when given a context of dialogue history along with conditions like persona, topic, and facts. Different from standard language understanding tasks, Chalamalasetti et al. (2023) presents a novel method to evaluate the ability of LLMs to act as ‘situational’ language understanding agents (Schlangen 2023). They do so by assigning roles to LLMs and generate dialogues resembling word games such as taboo, and test the language generating and instruction following abilities of LLMs based on the quality of game-play leading to successful target word prediction.
@@ 6 Conclusion
Although superlative performances have been reported on LLMs in recent times, recent work shows the performance gap between US English and other dialects of English. Our paper presents a first-of-its-kind evaluation of the dialect robustness of LLMs using their ability to predict target words in game-playing conversations. We use a dataset of target-word-masked conversations between US English speakers and those between Indian English speakers playing a game of taboo. We evaluate pre-trained and fine-tuned versions of one open-source and two closed-source models, on two tasks: target word prediction (TWP) and target word selection (TWS). Our results show that the LLMs indeed perform better for en-US as compared to en-IN on both tasks, with the average performance being higher by 12.66 and 17.4 on similarity and accuracy scores across all configurations. This shows that LLMs marginalise or discriminate against speakers of the Indian dialect. We also observe that pre-trained models report a degraded performance on conversations created using both rule-based (en-MV) and LLM-based (en-TR) transformations, as compared to their source conversations (en-US and en-IN respectively). However, fine-tuning on en-MV yields a greater improvement in the task performances, as compared to that on en-TR. This shows that the transformations that introduce dialectal information about a national variety help in improving the dialect robustness of LLMs more than the transformations that remove the said dialectal information. Finally, our error analysis demonstrates that, while most errors are mitigated by providing options for masked target word (TWS; in both pre-trained and fine-tuned variants), LLMs struggle to interpret target words based on the shared cultural context between speakers.
Our extension M-MD3 is a dataset for TWP and TWS based on MD3, consisting of four subsets: en-US, en-IN, en-MV, and en-TR. The dataset opens opportunities for future evaluations of dialect robustness using similar conversation-based tasks. Our evaluation methodology can also be applied to other existing dialogue and discourse datasets, to evaluate the ability of LLMs on properties other than dialect robustness.
@@
The original MD3 paper states that their dataset may overrepresent Western entities to some degree. Therefore, it is possible that Indian speakers faced difficulties with the terms. Having said that, the instances selected for our dataset are the ones where the Indian players guessed the word correctly. We have not performed a detailed qualitative analysis of these conversations, except for a cursory sanity check. We also assume that the dialect of English from each locale is homogeneous. Assuming that en-IN is the English spoken in every region of India is an unrealistic generalization of the diversity of dialects of English. In terms of model fine-tuning, our paper also does not cover the impact of quantization and different fine-tuning (including cross-dialect) techniques on the task.
@@
We use a publicly available dataset of conversations consisting of human players engaged in a game of taboo. The topics discussed in the dataset are fairly general, and are unlikely to cause distress. The error analysis was performed by one of the authors of the paper. The AI-transformed (en-TR) conversations may contain biased output, arising due to inherent properties of GPT-based models.

### v3

@@ Abstract
With an evergrowing number of LLMs reporting superlative performance for English, their ability to perform equitably for different dialects of English (i.e., dialect robustness) needs to be ascertained. Specifically, we use English language (US English or Indian English) conversations between humans who play the word-guessing game of ‘taboo‘. We formulate two evaluative tasks: target word prediction (TWP) (i.e., predict the masked target word in a conversation) and target word selection (TWS) (i.e., select the most likely masked target word in a conversation, from among a set of candidate words). Extending MD3, an existing dialectic dataset of taboo-playing conversations, we introduce M-MD3, a target-word-masked version of MD3 with the en-US and en-IN subsets. We create two subsets: en-MV (where en-US is transformed to include dialectal information) and en-TR (where dialectal information is removed from en-IN). We evaluate three multilingual LLMs–one open-source (Llama3) and two closed-source (GPT-4/3.5). LLMs perform significantly better for US English than Indian English for both TWP and TWS tasks, for all settings, exhibiting marginalisation against the Indian dialect of English. While GPT-based models perform the best, the comparatively smaller models work more equitably after fine-tuning. Our evaluation methodology exhibits a novel and reproducible way to examine attributes of language models using pre-existing dialogue datasets with language varieties. Dialect being an artifact of one’s culture, this paper demonstrates the gap in the performance of multilingual LLMs for communities that do not use a mainstream dialect.
Evaluating Dialect Robustness of Language Models via Conversation Understanding
Dipankar Srirag1 Nihar Sahoo2 Aditya Joshi1 1University of New South Wales, Sydney, Australia 2Indian Institute of Technology Bombay, India {d.srirag, aditya.joshi}@unsw.edu.au nihar@cse.iitb.ac.in
@@ 1 Introduction
refer to Indian English and US English respectively.Large language models (LLMs)111We use ‘language models’ and ‘large language models/LLMs’ interchangeably in this paper. based on Transformers Vaswani et al. (2017) are the state-of-the-art in natural language processing (NLP), often reporting superlative performance on several NLP tasks Zhao et al. (2023). These models predominantly use English language data in their pre-training corpus. However, being a widely spoken language, English takes multiple forms in different parts of the world. These forms, called dialects or national varieties of English, collectively constitute the World Englishes Bolton (2012). While research papers introducing LLMs report performance on English language datasets, recent works highlight the performance gap between US English and other dialects of English for several natural language processing tasks Joshi et al. (2024).
Our paper examines cultural considerations of evaluating LLMs through the prism of dialect robustness via conversation understanding. The choice of conversation understanding as a domain for evaluation emerges from the fact that dialectal features are most visible in free-flowing conversations Negro and Vietti (2006). Therefore, we investigate the research question:
@@
To address the research question, we use a pre-existing dataset– MD3 Eisenstein et al. (2023) that consists of manually transcribed dialogues between pairs of human participants where each pair speaks either Indian English or US English. The participants engage in a focused conversation: they play the word-guessing game based on the game of ‘Taboo’ Wikipedia (2023). In the game, a describer must get a guesser to identify a target word but must not use a set of related words known as restricted words while describing the target word. Using this dataset of dialectal dialogues, we introduce two tasks to evaluate the dialect-robustness of LLMs to understand conversations. They are: (a) Given an input conversation with the target word masked, can the LLM predict the target word? (referred to as target word prediction) (b) Given an input conversation with the target word masked along with a set of candidate target words, can the LLM select the correct target word? (referred to as target word selection). Our approach of masking the target word is similar to Dey and Desarkar (2023), who show that masked word prediction may correlate with automatic dialogue evaluation metrics. Figure 1 shows an example of the two tasks, where the language model predicts ‘Justin Bieber’ for target word prediction, and selects ‘microphone’ among the set of options for target word selection222We run experiments on both the tasks for both US and Indian English conversations. While the examples show expected output, the LLM may or may not produce the same in the case of our experiments. That is the crux of the evaluation.. For the two tasks, we extend MD3 to create a target-word-Masked Multi-Dialect Dataset of Dialogues (M-MD3)333M-MD3 dataset and the related code will be made publicly available at https://github.com/dipankarsrirag/eval-dialect-robust.. M-MD3 consists of (a) conversations between Indian English speakers (en-IN), and conversations between US English speakers (en-US), (b) en-US conversations transformed into en-IN using rule-based perturbations (en-MV), (c) en-IN with dialectal information removed (en-TR). We evaluate the performance of three SOTA large language models (LLMs), one open-source and two closed-source, employing zero-shot prompting on both pre-trained and fine-tuned models (where available). Our evaluation methodology derives from past work that evaluates LLMs by providing a set of task-specific examples Wang et al. (2023). Of particular relevance is the work by Chalamalasetti et al. (2023), who generate word game conversations using LLMs and evaluate their ability to predict the target word. The contributions of our work are:
@@
Our evaluation demonstrates a degraded performance in the case of Indian English as compared to US English for all models, supporting existing social disparities between US and Indian culture in the LLM representations Khandelwal et al. (2024).
@@
Since several LLMs have been deployed as publicly available dialogue agents444ChatGPT https://chat.openai.com/; Accessed on 9th April 2024., it is imperative that they can understand the conversations of users belonging to diverse English-speaking subgroups. In the case of our paper, this refers to dialectal variations, considering them as a proxy to culture. The rest of the paper is organized as follows. Section 2 introduces our evaluation methodology. The experiment setup and results are in Sections 3 and 4 respectively.
@@ 2 Methodology
We present our method step-by-step, with a detailed overview of our evaluation methodology described in Figure 2. We select two subsets available in MD3: en-IN and en-US, and filter out the conversations where the guesser could not identify the target word. We extend MD3 to include two additional sets of conversations—en-MV and en-TR, and mask the target words in all four subsets to create M-MD3. We ensure that the mask token always appears at the end of the conversation, warranting the use of auto-regressive models. This is done by pruning the conversation to the turn where the guesser utters the target word555Details on the masking method with examples are provided in Appendix A..
Transforming text in en-US to other dialectal English text has been explored for low-resource settings Held et al. (2023); Xiao et al. (2023); Liu et al. (2023). To evaluate the efficacy of synthetically transformed dialogues, we extend the dataset of dialectal dialogues to include two additional sets of conversations– en-MV and en-TR.
en-MV
We use Multi-VALUE Ziems et al. (2023) to transform en-US conversations into en-IN conversations. We call this set of conversations created by rule-based transformations en-MV.
en-TR
We prompt GPT-4 Turbo Preview(gpt-4; OpenAI et al. 2024) to remove dialectal information from en-IN. The resultant set of conversations is known as en-TR. The prompt666The forms are experimentally determined using a few test examples. used to generate such conversations is given below:
@@ 2.1 Extending MD3
The use of gpt-4 to transform en-IN conversations sometimes leads to the generation of conversation summaries rather than transformed conversations777More details with examples are discussed in Appendix B.. Due to the varying lengths of speaker turns, transforming en-US conversations using Multi-VALUE occasionally fails to output a result. Such failed transformations are excluded from both the subsets of transformed (en-MV, en-TR) conversations, leading to fewer conversations in en-MV and en-TR as compared to en-US and en-IN, respectively, as shown in Tables 1 and 2.
@@ 2.2 Analysis
We notice a higher number of average turns and words spoken in en-IN conversations compared to en-US conversations. This is due to the en-US speakers being more familiar with the target word compared to en-IN speakers, leading to shorter gameplay time Eisenstein et al. (2023). The trend is also carried over to the transformed conversations in en-MV (derived from en-US) and en-TR (derived from en-IN).
@@ 2.3 Task Definition
As shown in Figure 3, we mask the target word in the conversations from all four subsets. The target word occurs in the last dialogue turn of the conversation, which is spoken by the guesser888This always holds because of the way we process the conversations.. As a result, we formulate two tasks where the expected output is to fill the correct word at the masked position:
@@
We then use prompting on three LLMs to perform both tasks (TWP and TWS). As LLMs, we choose models that have been optimised to follow natural language instructions. In our case, the instruction is to either predict the masked target word or select a word from candidate words. Specifically, we use one open-source model, namely, Llama 3 70B Chat (llama-3; Grattafiori et al. 2024), and two closed-source models, namely, gpt-4 and GPT-3.5 Turbo 0125 (gpt-3.5; Ouyang et al. 2024).
@@ 3 Experiment Setup
We report the performance on pre-trained and fine-tuned versions of multilingual LLMs using zero-shot prompting. Fine-tuning is always done ‘in-dialect’ in our case, although there is no reason to believe that cross-dialect fine-tuning is not possible.
@@ 3.1 Model Parameters
Experiments on gpt-4 and gpt-3.5 are conducted using OpenAI’s API999OpenAI API https://platform.openai.com/docs/api-reference; Accessed on 18th April 2024.. gpt-3.5 is fine-tuned for 5 epochs, separately for every subset. We select top_p as 0.2 to restrict variability in output generation.
llama-3 is fine-tuned for 20 epochs, with a batch size of 16, Paged 8-bit AdamW Dettmers et al. (2022) as the optimiser and a learning rate of 2e-4. We use QLoRA adaptors, targeting all linear layers, as recommended by Dettmers et al. (2024). All experiments on llama-3 were performed using two A100 GPUs.
@@ 3.2 Metrics
We report our results on two metrics: accuracy and similarity. Accuracy is the proportion of conversations where the LLM generated the correct target word. This is a strict metric in that it requires the LLM to generate an exact match to the reference target word. In the case of TWP, the LLM will choose from all the words within its vocabulary, while in the case of TWS, the LLM will choose from the set of candidate target words. Therefore, it is trivial that the accuracy for TWS is expected to be higher than that for TWP. Accuracy metric penalizes models even if the generated target word partially matches with the reference target word in case of multi-word reference target as described in Section 2.2. As similarity, we report the cosine similarity between the Sentence-BERT embeddings Reimers and Gurevych (2019) of the reference target word and the generated target word. This allows for similar but inexact words generated by the LLM to be acceptable to the similarity score.
@@ 4.1 Quantitative Results
In the case of en-MV, the task performances are consistently lower compared to en-US. For example, in the case of gpt-3.5 and TWS, the similarity scores on the fine-tuned model are 80.8 for en-US and 71.5 for en-MV. This degraded performance shows that the rule-based transformation into en-IN from en-US reduced the understanding capacity of LLMs for the resulting conversations, further strengthening our hypothesis that LLMs perform well for US English dialects compared to any other varieties, similar to findings of Ryan et al. (2024).
@@ 4.2 Error Analysis
From Test set of each conversation subset, we randomly select 30 conversations that are mislabeled by gpt-4 and llama-3, and manually analyse errors among all model variants across all subsets of conversations. We summarise the six error categories101010Additonal examples for each error category are in Appendix C. in Table 4. The error types are:
@@ 5 Related Work
Research in dialect robustness stems from the need for language technologies to be equitable and not reinforce any negative sentiments against a specific linguistic subgroup Blodgett et al. (2020). LLMs perform poorly on several downstream tasks (such as the tasks in the GLUE benchmark) involving dialects other than mainstream US English Joshi et al. (2024); Faisal et al. (2024).
Similar to our work, the evaluation of language understanding ability of LLMs has been explored using typical conversation understanding tasks Chen et al. (2022) like conversation summarisation Gliwa et al. (2019); Chen et al. (2021), conversation completion Sai et al. (2020); Ueyama and Kano (2023), or NLU tasks Faisal et al. (2024). Other approaches involve conversation-based question-answering tasks that also evaluate the reasoning abilities of LLMs Sun et al. (2019); Qin et al. (2021). Tasks like mask-filling were used to evaluate LLM-generated responses, more specifically Dey and Desarkar (2023) do so by making RoBERTa predict masked keyword utterances when given a context of dialogue history along with conditions like persona, topic, and facts. Different from standard language understanding tasks, Chalamalasetti et al. (2023) presents a novel method to evaluate the ability of LLMs to act as ‘situational’ language understanding agents Schlangen (2023). They do so by assigning roles to LLMs and generate dialogues resembling word games such as taboo, and test the language generating and instruction following abilities of LLMs based on the quality of game-play leading to successful target word prediction.
@@ 6 Conclusion
Although superlative performances have been reported on LLMs in recent times, recent work shows the performance gap between US English and other dialects of English. Our paper presents a first-of-its-kind evaluation of the multilingual LLMs for their robustness to minority language varieties, using their ability to predict target words in game-playing conversations. We use a dataset of target-word-masked conversations between US English speakers and those between Indian English speakers playing a game of taboo. We evaluate pre-trained and fine-tuned versions of one open-source and two closed-source models, on two tasks: target word prediction (TWP) and target word selection (TWS). Our results show that the LLMs indeed perform better for en-US as compared to en-IN on both tasks, with the average performance being higher by 12.66 and 17.4 on similarity and accuracy scores across all configurations. This shows that the LLMs, although multilingual, marginalise or discriminate against speakers of the Indian dialect. We also observe that pre-trained models report a degraded performance on conversations created using both rule-based (en-MV) and LLM-based (en-TR) transformations, as compared to their source conversations (en-US and en-IN respectively). However, fine-tuning on en-MV yields a greater improvement in the task performances, as compared to that on en-TR. This shows that the transformations that introduce dialectal information about a national variety help in improving the dialect robustness of LLMs more than the transformations that remove the said dialectal information. Finally, our error analysis demonstrates that, while most errors are mitigated by providing options for masked target words (TWS; in both pre-trained and fine-tuned variants), multilingual LLMs struggle to interpret target words based on the shared cultural context between speakers.
Our extension M-MD3 is a dataset for TWP and TWS based on MD3, consisting of four subsets: en-US, en-IN, en-MV, and en-TR. The dataset opens opportunities for future evaluations of dialect robustness using similar conversation-based tasks. Our evaluation methodology can also be scaled up and applied to other existing dialogue and discourse datasets, to evaluate the ability of LLMs on properties other than dialect robustness.
@@
The original MD3 paper states that their dataset may be dominated by Western entities to some degree. Therefore, it is possible that Indian speakers faced difficulties with the terms. Having said that, the instances selected for our dataset are the ones where the Indian players guessed the word correctly. We have not performed a detailed qualitative analysis of these conversations, except for a cursory sanity check. We also assume that the dialect of English from each locale is homogeneous. Assuming that en-IN is the English spoken in every region of India is an unrealistic generalization of the diversity of dialects of English. In terms of model fine-tuning, our paper also does not cover the impact of quantization and different fine-tuning (including cross-dialect) techniques on the task.
@@
We use a publicly available dataset of conversations consisting of human players engaged in a game of taboo. The topics discussed in the dataset are fairly general and are unlikely to cause distress. The error analysis was performed by one of the authors of the paper. The AI-transformed (en-TR) conversations may contain biased output, arising due to inherent properties of GPT-based models.